# Personalized Predictions of Glioblastoma Infiltration: Mathematical Models, Physics-Informed Neural Networks and Multimodal Scans

### v2

@@ 1 Introduction
Glioblastoma (GBM) is a common and very aggressive form of brain tumor, with a median survival of 15 months (Stupp et al., 2014, 2005). The current standard of care for GBM includes immediate resection of the visible tumor after diagnosis, followed by radiotherapy and chemotherapy targeting residual tumor (Fernandes et al., 2017). Despite intensive treatment, the prognosis of GBM remains poor, as almost all GBMs recur. The high recurrence rate of GBM is attributed to its infiltrative nature. Instead of forming a clear boundary, the tumor invades into the surrounding normal-looking brain tissue (Souhami et al., 2004; Halperin et al., 1989) The infiltration pattern of tumor cells is heterogeneous and varies significantly among patients. Consequently, the standard clinical target volume (CTV) of radiotherapy, which extends 1.5-3 cm uniformly from the visible tumor, may be insufficient for optimal coverage and distribution (Stupp et al., 2005). A critical step towards enhancing radiotherapy treatment planning requires the estimation of tumor cell distributions in a data-driven and patient-specific manner (Lipková et al., 2019; Unkelbach et al., 2014; Rockne et al., 2010, 2015, 2019).
@@
Mathematical models of tumor growth are instrumental in predicting tumor progression and guiding individualized treatment decisions for patients (Baldock et al., 2013; Lipková et al., 2019; Butner et al., 2022; Ezhov et al., 2023; Chaudhuri et al., 2023; Subramanian et al., 2023; Lorenzo et al., 2023; Wu et al., 2022). The growth of GBM is often modeled using partial differential equations (PDEs) of reaction-diffusion type, which describe the spatio-temporal evolution of tumor cell density and captures the main tumor behavior: proliferation and infiltration (Harpold et al., 2007). More advanced models incorporate additional features including the mass effect (Hogea et al., 2007; Subramanian et al., 2019; Lipková et al., 2022), angiogenesis (Macklin et al., 2009; Saut et al., 2014; Yan et al., 2017), chemotaxis (Lu et al., 2022, 2023), and multiple species (Curtin et al., 2020; Yan et al., 2017). For comprehensive reviews on mathematical modeling of GBM, see Alfonso et al. (2017); Falco et al. (2021); Jørgensen et al. (2023).
@@
To apply these models in a clinical setting, it is imperative to estimate patient-specific parameters of the PDEs from clinical data, which is a challenging inverse problem. Besides being useful in solving the PDE to predict tumor cell density profiles, these patient-specific parameters are valuable on their own as biophysics-based features to improve patient stratification and survival prediction. For example, recent studies show that these biophysical features correlate better with survival time than volumetric features from images (Subramanian et al., 2023). Further, the ratio of proliferation rate to diffusion rate is key in evaluating the effectiveness of supramarginal resection for IDH-wildtype GBM patients (Tripathi et al., 2021).
@@
The inverse problem is commonly solved in the framework of PDE constrained optimization (Mang et al., 2012; Hogea et al., 2008; Scheufele et al., 2020, 2021; Subramanian et al., 2020b, a, 2023) or Bayesian inference (Menze et al., 2011; Lipková et al., 2019; Ezhov et al., 2019, 2022a). Machine learning approaches have also been explored to speed up the computations, including learning surrogate models for forward PDE systems (Ezhov et al., 2022a), learning mappings from observations to parameters directly (Pati et al., 2021; Ezhov et al., 2023; Martens et al., 2022), using dynamic mode decomposition to construct low-dimensional representations (Viguerie et al., 2022), and optimizing discrete loss frameworks that integrate the PDE and the data using mesh-based discretizations (Karnakov et al., 2022; Balcerak et al., 2023). Instead of the full reaction-diffusion model, its asymptotic properties can be exploited to obtain an approximate anisotropic Eikonal approximation, which simplifies parameter estimation and tumor cell density prediction, albeit with a less detailed model (Konukoglu et al., 2010a, b; Unkelbach et al., 2014).
@@
Physics-Informed Neural Networks (PINNs) have emerged as a novel method for solving inverse problems in a PDE constrained optimization framework (Karniadakis et al., 2021; Raissi et al., 2019; Jagtap et al., 2022a, b; Chen et al., 2020; Zheng et al., 2023; Huang et al., 2022). In this approach, PINNs represent the PDE solution as a neural network and embed both the data and the PDE, via a mesh-free approach using automatic differentiation, into the loss function. By minimizing the total loss, PINNs effectively solve the PDE and fit the data simultaneously, showcasing integration of mathematical models with data-driven learning processes. This is particularly beneficial in the realm of medical image analysis, where clinical measurements are often sparse and expensive. PINNs have been employed to combine the transcranial doppler data of blood flow velocity measurements and PDEs that govern the flow rate, pressure, and vessel area to estimate the brain hemodynamics (Sarabian et al., 2022). PINNs have also been used to calibrate the parameters in tracer-kinetic models with perfusion MR data for quantification of myocardial blood flow (van Herten et al., 2022). The potential for solving GBM inverse problems with PINNs was explored in Zhu (2022) using synthetic cell density data.
@@ 2.1.1 Tumor Model
We consider the Fisher-KPP PDE, a fundamental model that describes the spatio-temporal evolution of the normalized tumor cell density $u(\mathbf{x},t)$ in a 3D domain $\Omega$ consisting of white matter (WM) and grey matter (GM) regions (Harpold et al., 2007; Swanson et al., 2000). The PDE, with Neumann boundary conditions, is given by
@@
[…]
@@
where $D=D_{w}$ is the scalar diffusion coefficient in white matter, and $P(\mathbf{x})=P_{w}(\mathbf{x})+10P_{g}(\mathbf{x})$ defines the geometry of the brain. We use an atlas-guided approach to obtain the brain geometry and tissue percentages. In particular, we register the brain atlas used in Lipková et al. (2019) to patient T1w scans using rigid registration following Lipková et al. (2019) . While rigid registration works well in many cases, some errors may arise near the ventricles, especially when there is significant deformation of the brain due to the tumor. It remains an active area of research to design registration methods that robustly account for significant distortions due to tumor mass effect and post-surgical resection (Lipková et al., 2019). While we used a single brain atlas here, we could have used multiple atlases and chosen the atlas for registration that most closely matched the patient scans (Subramanian et al., 2023). We assume the initial condition is $u_{0}(\mathbf{x})=0.1\exp(-0.1|\mathbf{x}-{\mathbf{x}_{0}}|^{2})$, where ${\mathbf{x}_{0}}$ denotes the initial location of the tumor.
@@ 2.1.2 Diffuse Domain Method
A common way of enforcing the boundary condition using PINN requires sampling collocation points on the exact location of the boundary. However, the brain geometry is complex and the image resolution is limited, therefore an exact parameterization and the normal vector field of the boundary is not easily obtainable. In our work, the boundary condition is handled by the diffuse domain method (Lervåg and Lowengrub, 2015; Li et al., 2009), which embeds the complex brain region $\Omega$ into a cubic box $\Omega_{B}$ by using a phase field function $\phi$ that replaces the sharp boundary of the domain with a smooth transition layer, and the PDE is modified:
@@ 2.1.3 Scaling
The MRI image shows the thresholded tumor cell density at an unknown time $t_{\rm end}$. It is impossible to infer $(D,\rho,t_{\rm end})$ simultaneously from a single time point: doubling the time and halving $D$ and $\rho$ will give the same solution. In Konukoglu et al. (2010b); Ezhov et al. (2023, 2022b), time-independent parameters $\sqrt{Dt_{\rm end}}$ and $\sqrt{\rho t_{\rm end}}$ are considered, and they suffice to predict the tumor cell density and infiltration pattern at the unknown time of imaging. Additional information regarding the time scale would be needed to determine the precise values of $(D,\rho,t_{\rm end})$
In this work, we apply the idea from (Konukoglu et al., 2010b; Ezhov et al., 2023), but with a different scaling technique. Let $\bar{D}$ and $\bar{\rho}$ be characteristic values of the diffusion coefficient and proliferation rate. Then, we can define the characteristic velocity $\bar{v}=\sqrt{\bar{D}\bar{\rho}}$. Given a length scale $\bar{L}$, we can define a characteristic time $\bar{T}=\bar{L}/\bar{v}$.
@@ 2.1.4 Multimodal Imaging Model
We consider binary segmentations obtained from T1Gd, which highlights the tumor core, and FLAIR, which highlights the edema region in addition to the tumor core (Lê et al., 2016; Lipková et al., 2019; Subramanian et al., 2020b; Tunc et al., 2021; Menze et al., 2011; Ezhov et al., 2023; Konukoglu et al., 2010a). Let $y^{\rm T1Gd}$ be the indicator function of the tumor core and $y^{\rm FLAIR}$ be the indicator function of the tumor core and edema region. We assume that the $y^{\rm T1Gd}$ and $y^{\rm FLAIR}$ are related to the tumor cell density $u$ through thresholding at $u_{c}^{\rm T1Gd}$ and $u_{c}^{\rm FLAIR}$ respectively, that is, for $s\in\{\rm T1Gd,FLAIR\}$
@@
Following Lipková et al. (2019), we can also incorporate the 18-fluoride-fluoro-ethyl-tyrosine positron emission tomography (FET-PET) signal as additional data in the inference process. The FET-PET signal measures the metabolic activity of the tumor cells and is positively correlated to the tumor cellularity (Stockhammer et al., 2008; Hutterer et al., 2013). However, FET-PET (hereafter referred to as FET) is not routinely available in the clinic. Let $p$ be the normalized FET signal obtained by subtracting the patient-specific background activity, thresholding and normalizing to [0,1], and $y^{\rm FET}$ be the indicator function of the region with positive FET signal. We assume that, in the region with positive normalized FET signal, there is a linear relationship between $p$ and $u$:
@@
where the intercept $A$ and the slope $m$ are patient-specific and unknown. In Lipková et al. (2019), FET was found to be important to obtain accurate tumor cell density predictions. Here, we show that FET is not necessary in our method.
@@ 2.2 Workflow
For samples in $\bar{D}/\bar{\rho}\in[0.1,1]$ and $\bar{L}\in[10,90]$ (which generate spherical tumors with radii smaller than 120mm), we solve the simplified PDE to find the radii $R^{\rm FLAIR}_{sph}$ and $R^{\rm T1Gd}_{sph}$ at which the tumor cell density is equal to characteristic values of the thresholds, which are taken to be ${\bar{u}}_{c}^{\rm FLAIR}=0.35$ and ${\bar{u}}_{c}^{\rm T1Gd}=0.6$, respectively, motivated by results in Lipková et al. (2019). The pair of $\bar{D}/\bar{\rho}$ and $L$ that results in $R^{\rm FLAIR}_{sph}$ and $R^{\rm T1Gd}_{sph}$ that are closest to $R^{\rm FLAIR}_{\rm seg}$ and $R^{\rm T1Gd}_{\rm seg}$ will be used as the characteristic parameters. The details of this procedure are given in E. This procedure is very fast as the PDE is one-dimensional, independent of the patient data, and the quantities $R^{\rm FLAIR}_{sph}$ and $R^{\rm T1Gd}_{sph}$ can be precomputed and stored as a look-up table. In addition, we do not need to perform a very refined grid search, because the actual parameters will be fine-tuned by the PINN with patient data.
@@ 2.3 Loss functions
Because the solution is changing rapidly at early times and the tumor grows from the center, we sample the collocation points for PDE loss densely at early times and in the center of the tumor. This was found to work better than other spatial distributions of the collocation points, including adaptive sampling (Lu et al., 2021b) (see B for details).
@@
In the fine tuning stage, we also restrict the range of the trainable parameters: $\mu_{\mathcal{D}}\in[0.75,1.25]$, $\mu_{\mathcal{R}}\in[0.75,1.25]$, $u_{c}^{\rm FLAIR}\in[0.2,0.5]$, $u_{c}^{\rm T1Gd}\in[0.5,0.8]$, $m\in[0.8,1.2]$, $A\in[0,1]$. As discussed earlier, we expect $\mu_{\mathcal{D}}$ and $\mu_{\mathcal{R}}$ to be close to 1. The ranges for other parameters are based on the priors in (Lipková et al., 2019). The constraints are enforced by adding the quadratic penalty losses: if a parameter $\beta\in[\beta_{\min},\beta_{\max}]$, then
@@
where $\Theta=\{\mu_{\mathcal{D}},\mu_{\mathcal{R}},{\mathbf{x}_{0}},u_{c}^{\rm FLAIR},u_{c}^{\rm T1Gd},m,A\}$ are trainable parameters, and $w_{\rm SEG}$ and $w_{\rm FET}$ are the weights of the segmentation and FET loss respectively. If FET data is not used, then $w_{\rm FET}=0$. In this work we set $w_{\rm SEG}=1e-3$ and $w_{\rm FET}=1e-2$ if FET data is used, while all the other losses have unit weights. These weights are chosen such that the losses are of similar magnitude, such that the training is not dominated by one particular loss. Although we tried other choices of the weights, including adaptively changing them in time (Groenendijk et al., 2020; Maddu et al., 2022; Chen et al., 2018; Wang et al., 2021), we found empirically that this choice of weights works well for all our numerical experiments. See A for details on training the neural network and for examples of the dynamics of loss functions during training. See F for different choices of the data loss.

### v3

@@ 1 Introduction
Glioblastoma (GBM) is a common and very aggressive form of brain tumor, with a median survival of 15 months [69, 70]. The current standard of care for GBM includes immediate resection of the visible tumor after diagnosis, followed by radiotherapy and chemotherapy targeting residual tumor [17]. Despite intensive treatment, the prognosis of GBM remains poor, as almost all GBMs recur. The high recurrence rate of GBM is attributed to its infiltrative nature. Instead of forming a clear boundary, the tumor invades into the surrounding normal-looking brain tissue [67, 22] The infiltration pattern of tumor cells is heterogeneous and varies significantly among patients. Consequently, the standard clinical target volume (CTV) of radiotherapy, which extends 1.5-3 cm uniformly from the visible tumor, may be insufficient for optimal coverage and distribution [70]. A critical step towards enhancing radiotherapy treatment planning requires the estimation of tumor cell distributions in a data-driven and patient-specific manner [43, 78, 59, 61, 60].
@@
Mathematical models of tumor growth are instrumental in predicting tumor progression and guiding individualized treatment decisions for patients [4, 43, 5, 15, 6, 71, 45, 84]. The growth of GBM is often modeled using partial differential equations (PDEs) of reaction-diffusion type, which describe the spatio-temporal evolution of tumor cell density and captures the main tumor behavior: proliferation and infiltration [24]. More advanced models incorporate additional features including the mass effect [25, 72, 44], angiogenesis [50, 63, 86], chemotaxis [49, 48], and multiple species [10, 86]. For comprehensive reviews on mathematical modeling of GBM, see Alfonso et al. [2], Falco et al. [16], Jørgensen et al. [31].
@@
To apply these models in a clinical setting, it is imperative to estimate patient-specific parameters of the PDEs from clinical data, which is a challenging inverse problem. Besides being useful in solving the PDE to predict tumor cell density profiles, these patient-specific parameters are valuable on their own as biophysics-based features to improve patient stratification and survival prediction. For example, recent studies show that these biophysical features correlate better with survival time than volumetric features from images [71]. Further, the ratio of proliferation rate to diffusion rate is key in evaluating the effectiveness of supramarginal resection for IDH-wildtype GBM patients [76].
@@
The inverse problem is commonly solved in the framework of PDE constrained optimization [52, 26, 65, 64, 74, 73, 71] or Bayesian inference [54, 43, 12, 13]. Machine learning approaches have also been explored to speed up the computations, including learning surrogate models for forward PDE systems [13], learning mappings from observations to parameters directly [57, 15, 53], using dynamic mode decomposition to construct low-dimensional representations [80], and optimizing discrete loss frameworks that integrate the PDE and the data using mesh-based discretizations [32, 3]. Instead of the full reaction-diffusion model, its asymptotic properties can be exploited to obtain an approximate anisotropic Eikonal approximation, which simplifies parameter estimation and tumor cell density prediction, albeit with a less detailed model [35, 36, 78].
@@
Physics-Informed Neural Networks (PINNs) have emerged as a novel method for solving inverse problems in a PDE constrained optimization framework [33, 58, 29, 30, 8, 88, 27]. In this approach, PINNs represent the PDE solution as a neural network and embed both the data and the PDE, via a mesh-free approach using automatic differentiation, into the loss function. By minimizing the total loss, PINNs effectively solve the PDE and fit the data simultaneously, showcasing integration of mathematical models with data-driven learning processes. This is particularly beneficial in the realm of medical image analysis, where clinical measurements are often sparse and expensive. PINNs have been employed to combine the transcranial doppler data of blood flow velocity measurements and PDEs that govern the flow rate, pressure, and vessel area to estimate the brain hemodynamics [62]. PINNs have also been used to calibrate the parameters in tracer-kinetic models with perfusion MR data for quantification of myocardial blood flow [79]. The potential for solving GBM inverse problems with PINNs was explored in [89] using synthetic cell density data.
@@ 2.1.1 Tumor Model
We consider the Fisher-KPP PDE, a fundamental model that describes the spatio-temporal evolution of the normalized tumor cell density $u(\mathbf{x},t)$ in a 3D domain $\Omega$ consisting of white matter (WM) and grey matter (GM) regions [24, 75]. The PDE, with Neumann boundary conditions, is given by
@@
where $\rho$ [1/day] is the proliferation rate and $\mathbb{D}(\mathbf{x})$ [mm2/day] is the diffusion tensor, and $\mathbf{n}$ is the outward normal vector on the boundary $\partial\Omega$. We further assume that the diffusion is isotropic, and is a weighted sum of the diffusion coefficient in WM and GM, $\mathbb{D}(\mathbf{x})=D_{w}P_{w}(\mathbf{x})+D_{g}P_{g}(\mathbf{x})$, where $P_{w}$, $P_{g}$ are the percentages of white and grey matter, and $D_{w}$, $D_{g}$ [mm2/day] are the diffusion coefficient in white and grey matter, and $D_{w}=10D_{g}$ [54, 43, 75]. Therefore, the original PDE (1) can be written as
@@
where $D=D_{w}$ is the scalar diffusion coefficient in white matter, and $P(\mathbf{x})=P_{w}(\mathbf{x})+10P_{g}(\mathbf{x})$ defines the geometry of the brain. We use an atlas-guided approach to obtain the brain geometry and tissue percentages. In particular, we register the brain atlas used in [43] to patient T1w scans using rigid registration following [43] . While rigid registration works well in many cases, some errors may arise near the ventricles, especially when there is significant deformation of the brain due to the tumor. It remains an active area of research to design registration methods that robustly account for significant distortions due to tumor mass effect and post-surgical resection [43]. While we used a single brain atlas here, we could have used multiple atlases and chosen the atlas for registration that most closely matched the patient scans [71]. We assume the initial condition is $u_{0}(\mathbf{x})=0.1\exp(-0.1|\mathbf{x}-{\mathbf{x}_{0}}|^{2})$, where ${\mathbf{x}_{0}}$ denotes the initial location of the tumor.
@@ 2.1.2 Diffuse Domain Method
A common way of enforcing the boundary condition using PINN requires sampling collocation points on the exact location of the boundary. However, the brain geometry is complex and the image resolution is limited, therefore an exact parameterization and the normal vector field of the boundary is not easily obtainable. In our work, the boundary condition is handled by the diffuse domain method [40, 41], which embeds the complex brain region $\Omega$ into a cubic box $\Omega_{B}$ by using a phase field function $\phi$ that replaces the sharp boundary of the domain with a smooth transition layer, and the PDE is modified:
@@ 2.1.3 Scaling
The MRI image shows the thresholded tumor cell density at an unknown time $t_{\rm end}$. It is impossible to infer $(D,\rho,t_{\rm end})$ simultaneously from a single time point: doubling the time and halving $D$ and $\rho$ will give the same solution. In Konukoglu et al. [36], Ezhov et al. [15, 14], time-independent parameters $\sqrt{Dt_{\rm end}}$ and $\sqrt{\rho t_{\rm end}}$ are considered, and they suffice to predict the tumor cell density and infiltration pattern at the unknown time of imaging. Additional information regarding the time scale would be needed to determine the precise values of $(D,\rho,t_{\rm end})$
In this work, we apply the idea from [36, 15], but with a different scaling technique. Let $\bar{D}$ and $\bar{\rho}$ be characteristic values of the diffusion coefficient and proliferation rate. Then, we can define the characteristic velocity $\bar{v}=\sqrt{\bar{D}\bar{\rho}}$. Given a length scale $\bar{L}$, we can define a characteristic time $\bar{T}=\bar{L}/\bar{v}$.
@@ 2.1.4 Multimodal Imaging Model
We consider binary segmentations obtained from T1Gd, which highlights the tumor core, and FLAIR, which highlights the edema region in addition to the tumor core [39, 43, 74, 77, 54, 15, 35]. Let $y^{\rm T1Gd}$ be the indicator function of the tumor core and $y^{\rm FLAIR}$ be the indicator function of the tumor core and edema region. We assume that the $y^{\rm T1Gd}$ and $y^{\rm FLAIR}$ are related to the tumor cell density $u$ through thresholding at $u_{c}^{\rm T1Gd}$ and $u_{c}^{\rm FLAIR}$ respectively, that is, for $s\in\{\rm T1Gd,FLAIR\}$
@@
Following [43], we can also incorporate the 18-fluoride-fluoro-ethyl-tyrosine positron emission tomography (FET-PET) signal as additional data in the inference process. The FET-PET signal measures the metabolic activity of the tumor cells and is positively correlated to the tumor cellularity [68, 28]. However, FET-PET (hereafter referred to as FET) is not routinely available in the clinic. Let $p$ be the normalized FET signal obtained by subtracting the patient-specific background activity, thresholding and normalizing to [0,1], and $y^{\rm FET}$ be the indicator function of the region with positive FET signal. We assume that, in the region with positive normalized FET signal, there is a linear relationship between $p$ and $u$:
@@
where the intercept $A$ and the slope $m$ are patient-specific and unknown. In Lipková et al. [43], FET was found to be important to obtain accurate tumor cell density predictions. Here, we show that FET is not necessary in our method.
@@ 2.2 Workflow
For samples in $\bar{D}/\bar{\rho}\in[0.1,1]$ and $\bar{L}\in[10,90]$ (which generate spherical tumors with radii smaller than 120mm), we solve the simplified PDE to find the radii $R^{\rm FLAIR}_{sph}$ and $R^{\rm T1Gd}_{sph}$ at which the tumor cell density is equal to characteristic values of the thresholds, which are taken to be ${\bar{u}}_{c}^{\rm FLAIR}=0.35$ and ${\bar{u}}_{c}^{\rm T1Gd}=0.6$, respectively, motivated by results in [43]. The pair of $\bar{D}/\bar{\rho}$ and $L$ that results in $R^{\rm FLAIR}_{sph}$ and $R^{\rm T1Gd}_{sph}$ that are closest to $R^{\rm FLAIR}_{\rm seg}$ and $R^{\rm T1Gd}_{\rm seg}$ will be used as the characteristic parameters. The details of this procedure are given in E. This procedure is very fast as the PDE is one-dimensional, independent of the patient data, and the quantities $R^{\rm FLAIR}_{sph}$ and $R^{\rm T1Gd}_{sph}$ can be precomputed and stored as a look-up table. In addition, we do not need to perform a very refined grid search, because the actual parameters will be fine-tuned by the PINN with patient data.
@@ 2.3 Loss functions
Because the solution is changing rapidly at early times and the tumor grows from the center, we sample the collocation points for PDE loss densely at early times and in the center of the tumor. This was found to work better than other spatial distributions of the collocation points, including adaptive sampling [47] (see B for details).
@@
In the fine tuning stage, we also restrict the range of the trainable parameters: $\mu_{\mathcal{D}}\in[0.75,1.25]$, $\mu_{\mathcal{R}}\in[0.75,1.25]$, $u_{c}^{\rm FLAIR}\in[0.2,0.5]$, $u_{c}^{\rm T1Gd}\in[0.5,0.8]$, $m\in[0.8,1.2]$, $A\in[0,1]$. As discussed earlier, we expect $\mu_{\mathcal{D}}$ and $\mu_{\mathcal{R}}$ to be close to 1. The ranges for other parameters are based on the priors in [43]. The constraints are enforced by adding the quadratic penalty losses: if a parameter $\beta\in[\beta_{\min},\beta_{\max}]$, then
@@
where $\Theta=\{\mu_{\mathcal{D}},\mu_{\mathcal{R}},{\mathbf{x}_{0}},u_{c}^{\rm FLAIR},u_{c}^{\rm T1Gd},m,A\}$ are trainable parameters, and $w_{\rm SEG}$ and $w_{\rm FET}$ are the weights of the segmentation and FET loss respectively. If FET data is not used, then $w_{\rm FET}=0$. In this work we set $w_{\rm SEG}=1e-3$ and $w_{\rm FET}=1e-2$ if FET data is used, while all the other losses have unit weights. These weights are chosen such that the losses are of similar magnitude, such that the training is not dominated by one particular loss. Although we tried other choices of the weights, including adaptively changing them in time [21, 51, 9, 82], we found empirically that this choice of weights works well for all our numerical experiments. See A for details on training the neural network and for examples of the dynamics of loss functions during training. See F for different choices of the data loss.